# The Complexity of Resilience Problems via Valued Constraint Satisfaction

### v5

@@ 4. Expressive Power
Now suppose that $R^{\Delta}=\operatorname{Feas}(S^{\Gamma})$ for some $S\in\tau$. Let $(\phi,u)$ be an instance of $\operatorname{VCSP}(\Delta)$, i.e., $\phi(x_{1},\dots,x_{n})=\sum_{i=1}^{\ell}\phi_{i}+\sum_{j=1}^{k}\psi_{j}$ where $\psi_{j}$, $j\in\{1,\dots,k\}$ are all the atomic expressions in $\phi$ that involve $R$. If $R^{\Delta}=R_{\emptyset}$, then the statement follows from the reduction for $R_{\emptyset}$. Therefore, suppose that this not the case. Since $\tau$ is finite and $\operatorname{Aut}(\Gamma)$ is oligomorphic, we may assume without loss of generality that all valued relations attain only non-negative values; otherwise we shift the values, which by the previous case does not affect the complexity up to polynomial-time reductions. Let $w$ be the maximum finite weight assigned by $S$. Note that there are only finitely many values that the $\ell$ atoms $\phi_{i}$ may take and therefore only finitely many values that $\sum_{i=1}^{\ell}\phi_{i}$ may take. Let $v$ be the smallest of these values such that $v>u$ and let $d=v-u$; if $v$ does not exist, let $d=1$. To simplify the notation, set $t=\lceil(kw)/d\rceil+1$. Let $\psi_{j}^{\prime}$ be the $\tau$-expression resulting from $\psi_{j}$ by replacing the symbol $R$ by the symbol $S$. Let $\phi^{\prime}$ be the $\tau$-expression obtained from $\phi$ by replacing each atom $\phi_{i}$ with $t$ copies of it and replacing every atom $\psi_{j}$ by $\psi_{j}^{\prime}$. Let $(\phi^{\prime},tu+kw)$ be the resulting instance of $\operatorname{VCSP}(\Gamma)$; note that it can be computed in polynomial time.
@@
Finally, we consider the case that $R^{\Delta}=\operatorname{Opt}(S^{\Gamma})$ for some relation symbol $S\in\tau$. Similarly to the previous case, we may assume without loss of generality that the minimum weight of all valued relations in $\Delta$ equals 00; otherwise, we subtract the smallest weight assigned to a tuple by some valued relation in $\Delta$. We may also assume that $S^{\Gamma}$ takes finite positive values, because otherwise $\operatorname{Opt}(S^{\Gamma})=S^{\Gamma}$ and the statement is trivial. Let $m$ be the smallest positive weight assigned by $S^{\Gamma}$ and let $M$ be the largest finite weight assigned by any valued relation of $\Gamma$ (again we use that $\tau$ is finite and that $\operatorname{Aut}(\Gamma)$ is oligomorphic). Let $(\phi,u)$, where $\phi(x_{1},\dots,x_{n})=\sum_{i=1}^{k}\phi_{i}$, be an instance of $\operatorname{VCSP}(\Delta)$. For $i\in\{1,\dots,k\}$, if $\phi_{i}$ involves the symbol $R$, then replace it by $k\cdot\lceil M/m\rceil+1$ copies and replace $R$ by $S$. Let $\phi^{\prime}$ be the resulting $\tau$-expression. We claim that $a\in C^{n}$ is a solution to the instance $(\phi^{\prime},\min(kM,u))$ of $\operatorname{VCSP}(\Gamma)$ if and only if it is the solution to $(\phi,u)$.
If $a\in C^{n}$ is such that $\phi(a)\leq u$ then for every $i\in\{1,\dots,k\}$ such that $\phi_{i}$ involves $R$ we have $\phi_{i}(a)=0$. In particular, the minimal value attained by $S^{\Gamma}$ equals 00 by our assumption, and hence $\phi^{\prime}(a)=\phi(a)\leq u$ and hence $\phi^{\prime}(a)\leq\min(kM,u)$. Now suppose that $\phi(a)>u$. Then $\phi^{\prime}(a)>u\geq\min(kM,u)$ or there exists an $i\in\{1,\dots,k\}$ such that $\phi_{i}(a)=\infty$. If $\phi_{i}$ does not involve the symbol $R$, then $\phi^{\prime}(a)=\infty$ as well. If $\phi_{i}$ involves the symbol $R$, then $\phi^{\prime}(a)\geq(k\cdot\lceil M/m\rceil+1)m>kM$. In any case, $\phi^{\prime}(a)>\min(kM,u)$. Since $\phi^{\prime}$ can be computed from $\phi$ in polynomial time, this concludes the proof. ∎
@@
We revisit the countably infinite valued structure $\Gamma_{\text{LCC}}$ from Example 3.2. Recall that $\operatorname{VCSP}(\Gamma_{\text{LCC}})$ is the least correlation clustering problem with partial information and that $\operatorname{Aut}(\Gamma_{\text{LCC}})$ is oligomorphic. Let $\Gamma_{\text{EC}}$ be the relational structure with the same domain as $\Gamma_{\text{LCC}}$ and the relation $R:=\{(x,y,z)\mid(x=y\wedge y\neq z)\vee(x\neq y\wedge y=z)\}$ (attaining values 00 and $\infty$). Note that
@@
Note that we can replace $N(x,z)+N(x,z)$ in the definition of $R$ by $\operatorname{Opt}(N)(x,z)$ and that $\operatorname{Opt}(N)$ is equal to the classical relation $\neq$ (attaining values 00 and $\infty$). This shows that even $\operatorname{VCSP}(\mathbb{N};E,\neq)$ is NP-hard.
@@ 5. Hardness from pp-Constructions
The name ‘pp-power’ comes from ‘primitive positive power’, since for relational structures expressibility is captured by primitive positive formulas. The following proposition shows that the VCSP of a pp-power reduces to the VCSP of the original structure.
@@
Let $d$ be the dimension of the pp-power and let $\tau$ be the signature of $\Gamma$. By Remark 4.5, $\operatorname{Aut}(\Gamma)\subseteq\operatorname{Aut}(\Delta)$ and thus $\operatorname{Aut}(\Delta)$ is oligomorphic. By Lemma 4.6, we may suppose that for every valued relation $R$ of arity $k$ of $\Delta$ the valued relation $S\in\langle\Gamma\rangle$ of arity $dk$ from the definition of pp-powers equals $S^{\Gamma}$ for some $S\in\tau$. Let $(\phi,u)$ be an instance of $\operatorname{VCSP}(\Delta)$. For each variable $x$ of $\phi$ we introduce $d$ new variables $x_{1},\dots,x_{d}$. For each summand $R(y^{1},\dots,y^{k})$ of $\phi$ we introduce a summand $S(y^{1}_{1},\dots,y^{1}_{d},\dots,y^{k}_{1},\dots,y^{k}_{d})$; let $\psi$ be the resulting $\tau$-expression. It is now straightforward to verify that $(\phi,u)$ has a solution with respect to $\Delta$ if and only if $(\psi,u)$ has a solution with respect to $\Gamma$. ∎
If $C$ and $D$ are sets, then we equip the space $C^{D}$ of functions from $D$ to $C$ with the topology of pointwise convergence, where $C$ is taken to be discrete. In this topology, a basis of open sets is given by
@@
for $a\in D^{k}$ and $b\in C^{k}$ for some $k\in{\mathbb{N}}$, and $f$ is applied componentwise. For any topological space $T$, we denote by $B(T)$ the Borel $\sigma$-algebra on $T$, i.e., the smallest subset of the powerset ${\mathcal{P}}(T)$ which contains all open sets and is closed under countable intersection and complement. We write $[0,1]$ for the set $\{x\in{\mathbb{R}}\mid 0\leq x\leq 1\}$.
@@
$(C^{D},B(C^{D}),\omega\colon B(C^{D})\to[0,1]),$
that is, $\omega(C^{D})=1$ and $\omega$ is countably additive: if $A_{1},A_{2},\dots\in B(C^{D})$ are disjoint, then
@@
If $f\in C^{D}$, we often write $\omega(f)$ instead of $\omega(\{f\})$. Note that $\{f\}\in B(C^{D})$ for every $f$. The set $[0,1]$ carries the topology inherited from the standard topology on ${\mathbb{R}}$. We also view ${}{\mathbb{R}}\cup\{\infty\}$ as a topological space with a basis of open sets given by all open intervals $(a,b)$ for $a,b\in{\mathbb{R}}$, $a<b$ and additionally all sets of the form $\{x\in{\mathbb{R}}\mid x>a\}\cup\{\infty\}$.
A (real-valued) random variable is a measurable function $X\colon T\to{\mathbb{R}}\cup\{\infty\}$, i.e., pre-images of elements of $B({\mathbb{R}}\cup\{\infty\})$ under $X$ are in $B(T)$. If $X$ is a real-valued random variable, then the expected value of $X$ (with respect to a probability distribution $\omega$) is denoted by $E_{\omega}[X]$ and is defined via the Lebesgue integral
[…]
@@
Let $C$ and $D$ be sets. In the rest of the article, we will work exclusively on a topological space $C^{D}$ of maps $f\colon D\rightarrow C$ and the special case ${\mathscr{O}}_{C}^{(\ell)}$ for some $\ell\in\mathbb{N}$ (i.e., $D=C^{\ell}$). Note that if $C$ and $D$ are infinite, then these spaces are uncountable and hence there are probability distributions $\omega$ such that $\omega(A)=0$ for every $1$-element set $A$. Therefore, in these cases, $E_{\omega}[X]$ for a random variable $X$ might not be expressible as a sum.
@@
Let $\Gamma$ and $\Delta$ be valued $\tau$-structures with domains $C$ and $D$, respectively. A fractional homomorphism from $\Delta$ to $\Gamma$ is a fractional map from $D$ to $C$ such that for every $R\in\tau$ of arity $k$ and every tuple $a\in D^{k}$ it holds for the random variable $X\colon C^{D}\rightarrow\mathbb{R}\cup\{\infty\}$ given by
@@
that $E_{\omega}[X]$ exists and that
@@
Let $C$ and $D$ be sets, $a\in D^{k}$, $R\in{\mathscr{R}}_{C}^{(k)}$. Let $X\colon C^{D}\rightarrow\mathbb{R}\cup\{\infty\}$ be the random variable given by
$f\mapsto R(f(a)).$
If $\operatorname{Aut}(C;R)$ is oligomorphic, then $E_{\omega}[X]$ exists and $E_{\omega}[X]>-\infty$.
@@
[…]
@@
Let $\Gamma_{1}$, $\Gamma_{2}$, $\Gamma_{3}$ be countable valued $\tau$-structures such that there exists a fractional homomorphism $\omega_{1}$ from $\Gamma_{1}$ to $\Gamma_{2}$ and a fractional homomorphism $\omega_{2}$ from $\Gamma_{2}$ to $\Gamma_{3}$. Then there exists a fractional homomorphism $\omega_{3}:=\omega_{2}\circ\omega_{1}$ from $\Gamma_{1}$ to $\Gamma_{3}$.

### v6

@@ 4. Expressive Power
Now suppose that $R^{\Delta}=\operatorname{Feas}(S^{\Gamma})$ for some $S\in\tau$. Let $(\phi,u)$ be an instance of $\operatorname{VCSP}(\Delta)$, i.e., $\phi(x_{1},\dots,x_{n})=\sum_{i=1}^{\ell}\phi_{i}+\sum_{j=1}^{k}\psi_{j}$ where $\psi_{j}$, $j\in\{1,\dots,k\}$ are all the atomic expressions in $\phi$ that involve $R$. If $R^{\Delta}=R_{\emptyset}$, then the statement follows from the reduction for $R_{\emptyset}$. Therefore, suppose that this not the case. Since $\tau$ is finite and $\operatorname{Aut}(\Gamma)$ is oligomorphic, we may assume without loss of generality that all valued relations attain only non-negative values; otherwise we shift the values, which by the previous case does not affect the complexity up to polynomial-time reductions. Let $w$ be the maximum finite weight assigned by $S$. Note that there are only finitely many values that the $\ell$ atoms $\phi_{i}$ may take and therefore only finitely many values that $\sum_{i=1}^{\ell}\phi_{i}$ may take. Let $v$ be the smallest of these values such that $v>u$ and let d=vud=v-u; if $v$ does not exist, let $d=1$. To simplify the notation, set $t=\lceil(kw)/d\rceil+1$. Let $\psi_{j}^{\prime}$ be the atomic expression resulting from $\psi_{j}$ by replacing the symbol $R$ by the symbol $S$. Let $\phi^{\prime}$ be the $\tau$-expression obtained from $\phi$ by replacing each atom $\phi_{i}$ with $t$ copies of it and replacing every atom $\psi_{j}$ by $\psi_{j}^{\prime}$. Let $(\phi^{\prime},tu+kw)$ be the resulting instance of $\operatorname{VCSP}(\Gamma)$; note that it can be computed in polynomial time.
@@
Finally, we consider the case that $R^{\Delta}=\operatorname{Opt}(S^{\Gamma})$ for some relation symbol $S\in\tau$. Similarly to the previous case, we may assume without loss of generality that the minimum weight of all valued relations in $\Delta$ equals 0; otherwise, we subtract the smallest weight assigned to a tuple by some valued relation in $\Delta$. We may also assume that $S^{\Gamma}$ takes finite positive values, because otherwise $\operatorname{Opt}(S^{\Gamma})=S^{\Gamma}$ and the statement is trivial. Let $m$ be the smallest positive weight assigned by $S^{\Gamma}$ and let $M$ be the largest finite weight assigned by any valued relation of $\Gamma$ (again we use that $\tau$ is finite and that $\operatorname{Aut}(\Gamma)$ is oligomorphic). Let $(\phi,u)$, where $\phi(x_{1},\dots,x_{n})=\sum_{i=1}^{k}\phi_{i}$, be an instance of $\operatorname{VCSP}(\Delta)$. For $i\in\{1,\dots,k\}$, if $\phi_{i}$ involves the symbol $R$, then replace it by $k\cdot\lceil M/m\rceil+1$ copies and replace $R$ by $S$. Let $\phi^{\prime}$ be the resulting $\tau$-expression. We claim that $a\in C^{n}$ is a solution to the instance $(\phi^{\prime},\min(kM,u))$ of $\operatorname{VCSP}(\Gamma)$ if and only if it is the solution to $(\phi,u)$.
If $a\in C^{n}$ is such that $\phi(a)\leq u$ then for every $i\in\{1,\dots,k\}$ such that $\phi_{i}$ involves $R$ we have $\phi_{i}(a)=0$. In particular, the minimal value attained by $S^{\Gamma}$ equals 0 by our assumption, and hence $\phi^{\prime}(a)=\phi(a)\leq u$ and hence $\phi^{\prime}(a)\leq\min(kM,u)$. Now suppose that $\phi(a)>u$. Then $\phi^{\prime}(a)>u\geq\min(kM,u)$ or there exists an $i\in\{1,\dots,k\}$ such that $\phi_{i}(a)=\infty$. If $\phi_{i}$ does not involve the symbol $R$, then $\phi^{\prime}(a)=\infty$ as well. If $\phi_{i}$ involves the symbol $R$, then $\phi^{\prime}(a)\geq(k\cdot\lceil M/m\rceil+1)m>kM$. In any case, $\phi^{\prime}(a)>\min(kM,u)$. Since $\phi^{\prime}$ can be computed from $\phi$ in polynomial time, this concludes the proof. ∎
@@
We revisit the countably infinite valued structure $\Gamma_{\text{LCC}}$ from Example 3.2. Recall that $\operatorname{VCSP}(\Gamma_{\text{LCC}})$ is the least correlation clustering problem with partial information and that $\operatorname{Aut}(\Gamma_{\text{LCC}})$ is oligomorphic. Let $\Gamma_{\text{EC}}$ be the relational structure with the same domain as $\Gamma_{\text{LCC}}$ and the relation $R:=\{(x,y,z)\mid(x=y\wedge y\neq z)\vee(x\neq y\wedge y=z)\}$ (attaining values 0 and $\infty$). Note that
@@
Note that we can replace $N(x,z)+N(x,z)$ in the definition of $R$ by $\operatorname{Opt}(N)(x,z)$ and that $\operatorname{Opt}(N)$ is equal to the classical relation $\neq$ (attaining values 0 and $\infty$). This shows that even $\operatorname{VCSP}(\mathbb{N};E,\neq)$ is NP-hard.
@@ 5. Hardness from pp-Constructions
The name ‘pp-power’ comes from ‘primitive positive power’, since for relational structures pp-expressibility is captured by primitive positive formulas. The following proposition shows that the VCSP of a pp-power reduces to the VCSP of the original structure.
@@
Let $d$ be the dimension of the pp-power and let $\tau$ be the signature of $\Gamma$. By Remark 4.4, $\operatorname{Aut}(\Gamma)\subseteq\operatorname{Aut}(\Delta)$ and thus $\operatorname{Aut}(\Delta)$ is oligomorphic. By Lemma 4.5, we may suppose that for every valued relation $R$ of arity $k$ of $\Delta$ the valued relation $S\in\langle\Gamma\rangle$ of arity $dk$ from the definition of a pp-power equals $S^{\Gamma}$ for some $S\in\tau$. Let $(\phi,u)$ be an instance of $\operatorname{VCSP}(\Delta)$. For each variable $x$ of $\phi$ we introduce $d$ new variables $x_{1},\dots,x_{d}$. For each summand $R(y^{1},\dots,y^{k})$ of $\phi$ we introduce a summand $S(y^{1}_{1},\dots,y^{1}_{d},\dots,y^{k}_{1},\dots,y^{k}_{d})$; let $\psi$ be the resulting $\tau$-expression. It is now straightforward to verify that $(\phi,u)$ has a solution with respect to $\Delta$ if and only if $(\psi,u)$ has a solution with respect to $\Gamma$. ∎
Heading towards the definition of pp-constructions, we introduce the important notion of fractional maps. To establish this notion and its basic properties, we need to work with notions from topology and measure theory, which we carefully introduce below at the required level of generality. If $C$ and $D$ are sets, then we equip the space $C^{D}$ of functions from $D$ to $C$ with the topology of pointwise convergence, i.e., the product topology on the space $C^{D}$ where $C$ is a discrete topological space. In this topology, a basis of open sets is given by
@@
for $a\in D^{k}$ and $b\in C^{k}$ for some $k\in{\mathbb{N}}$, and $f$ is applied componentwise. For any topological space $T$, we denote by $\mathcal{B}(T)$ the Borel $\sigma$-algebra on $T$, i.e., the smallest subset of the powerset ${\mathcal{P}}(T)$ which contains all open sets and is closed under countable intersection and complement. We write $[0,1]$ for the set $\{x\in{\mathbb{R}}\mid 0\leq x\leq 1\}$.
@@
$(C^{D},\mathcal{B}(C^{D}),\omega\colon\mathcal{B}(C^{D})\to[0,1]),$
that is, $\omega(C^{D})=1$ and $\omega$ is countably additive: if $A_{1},A_{2},\dots\in\mathcal{B}(C^{D})$ are disjoint, then
@@
Note that every probability distribution is a measure, i.e., it is a countably additive map $\omega$ from a $\sigma$-algebra to non-negative real numbers such that $\omega(\emptyset)=0$. If $f\in C^{D}$, we often write $\omega(f)$ instead of $\omega(\{f\})$. Note that $\{f\}\in\mathcal{B}(C^{D})$ for every $f$. The set $[0,1]$ carries the topology inherited from the standard topology on $\mathbb{R}$}. We also view $\mathbb{R}$$\cup\{\infty\}$ as a topological space with a basis of open sets given by all open intervals $(a,b)$ for $\in{\mathbb{R}}, a<ba<b$ and additionally all sets of the form $\{x\in{\mathbb{R}}\mid x>a\}\cup\{\infty\}$.
A function $f\colon S\to T$ between two topological spaces is called (Borel-) measurable if $f^{-1}(T^{\prime})\in\mathcal{B}(S)$ for every $T^{\prime}\in\mathcal{B}(T)$. A (real-valued) random variable is a measurable function $X\colon T\to{\mathbb{R}}\cup\{\infty\}$. If $X$ is a real-valued random variable, then the expected value of $X$ (with respect to a probability distribution $\omega$) is denoted by $E_{\omega}[X]$ and is defined via the Lebesgue integral
$E_{\omega}[X]:=\int_{T}Xd\omega.$
Recall that the Lebesgue integral $\int_{T}Xd\omega$ need not exist, in which case $E_{\omega}[X]$ is undefined; otherwise, the integral equals a real number, $\infty$, or -$\infty$. The definition and some properties of the Lebesgue integral, specialized to our setting, can be found in Appendix A. Also recall that the expected value is
@@
Let $C$ and $D$ be sets. In the rest of the article, we will work exclusively on a topological space $C^{D}$ of maps $f\colon D\rightarrow C$ and the special case where $D=C^{\ell}$ for some $\ell\in\mathbb{N}$. Note that if $C$ and $D$ are infinite, then these spaces are uncountable and hence there are probability distributions $\omega$ such that $\omega(A)=0$ for every $1$-element set $A$. Therefore, in these cases, $E_{\omega}[X]$ for a random variable $X$ might not be expressible as a sum.
@@
Let $\Gamma$ and $\Delta$ be valued $\tau$-structures with domains $C$ and $D$, respectively. A fractional homomorphism from $\Delta$ to $\Gamma$ is a fractional map from $D$ to $C$ such that for every $R\in\tau$ of arity $k$ and every tuple $a\in D^{k}$, the expected value $E_{\omega}[X]$ of the random variable $X\colon C^{D}\rightarrow\mathbb{R}\cup\{\infty\}$ given by
@@
exists and satisfies
@@
Note that fractional homomorphisms between valued structures are a natural generalization of homomorphisms between relational structures: if $\Delta$ and $\Gamma$ are 0-$\infty$-valued and $\omega(\{f\})=1$ for some $f\colon D\to C$, then the definition requires that $R^{\Gamma}(f(a))\leq R^{\Delta}(a)$ for every $a$, equivalently, whenever $R^{\Delta}(a)=0$, then $R^{\Gamma}(f(a))=0$, which is equivalent to $f$ being a homomorphism.
@@
Let $C$ and $D$ be sets, $a\in D^{k}$, and $R\in{\mathscr{R}}_{C}^{(k)}$. If $X\colon C^{D}\rightarrow\mathbb{R}\cup\{\infty\}$ is the random variable given by
$f\mapsto R(f(a))$
and $\operatorname{Aut}(C;R)$ is oligomorphic, then $E_{\omega}[X]$ exists and $E_{\omega}[X]>$-$\infty$.
@@
It is enough to show that $\int_{C^{D}}X^{-}d\omega\neq\infty$. Since $\operatorname{Aut}(C;R)$ is oligomorphic, there are only finitely many orbits of $k$-tuples in $\operatorname{Aut}(C;R)$. Let $O_{1},\dots,O_{m}$ be all orbits of $k$-tuples of $\operatorname{Aut}(C;R)$ on which $R$ is negative. For every $i\in\{1,\dots,m\}$, let $b_{i}\in O_{i}$. Then we obtain (see (27) in Appendix A for a detailed derivation of the first equality)
$\int_{C^{D}}X^{-}d\omega$ $=\sum_{b\in C^{k},R(b)<0}-R(b)\omega({\mathscr{S}}_{a,b})$
@@
Let $T_{1}$ and $T_{2}$ be topological spaces. Then the product ${\mathcal{B}}(T_{1})\otimes{\mathcal{B}}(T_{2})$ is the smallest $\sigma$-algebra such that for $i\in\{1,2\}$, the projection $\pi_{i}\colon T_{1}\times T_{2}\to T_{i}$ given by $\pi_{i}(U_{1}\times U_{2}):=U_{i}$ is measurable. For $i\in\{1,2\}$, let $\omega_{i}$ be a probability distribution on $T_{i}$. Then the probability distribution $\omega_{1}\otimes\omega_{2}$ on $\mathcal{B}(T_{1})\otimes\mathcal{B}(T_{2})$ is the unique probability distribution that satisfies for all $U_{1}\in\mathcal{B}(T_{1})$ and $U_{2}\in\mathcal{B}(T_{2})$ that
$\omega_{1}\otimes\omega_{2}(U_{1}\times U_{2})=\omega_{1}(U_{1})\cdot\omega_{2}(U_{2});$
the existence and uniqueness of this measure is folklore, see, e.g., (Klenke, 2020, Theorem 14.14). A topological space $S$ is called second-countable if it has a countable basis of open sets; note that if $C$ and $D$ are countable, then the product space $C^{D}$ is second-countable. We provide the following statement with a proof for the convenience of the reader.
Lemma 5.6 (see, e.g., (Bogachev, 2007, Volume II, Lemma 6.4.2) or (Gheysens, 2017, Lemma D.1.1)).
Let $S$ and $T$ be second-countable topological spaces. Then ${\mathcal{B}}(S)\otimes{\mathcal{B}}(T)={\mathcal{B}}(S\times T)$.
Proof.
The inclusion $\subseteq$ holds without the assumption of second-countability. To see this, note that ${\mathcal{B}}(S)\otimes{\mathcal{B}}(T)$ is the coarsest $\sigma$-algebra such that the projections $\pi_{i}$ are measurable whereas ${\mathcal{B}}(S\times T)$ is the Borel $\sigma$-algebra of the coarsest topology such that the projections are continuous. It is easy to see that if a map is continous with respect to a topology, then it is measurable over the Borel $\sigma$-algebra generated by this topology, hence we must have ${\mathcal{B}}(S)\otimes{\mathcal{B}}(T)\subseteq{\mathcal{B}}(S\times T)$.
To see the reverse inclusion, note that the product $S\times T$ is also second-countable with a countable basis $\mathscr{B}$ of open sets. Each element of $\mathscr{B}$ trivially belongs to ${\mathcal{B}}(S\times T)$: this is because the projections $\pi_{i}$ are measurable with respect to $\mathcal{B}(S)\otimes{\mathcal{B}}(T)$ and hence for every $S^{\prime}\times T^{\prime}\in\mathscr{B}$ the sets $S^{\prime}\times T$ and $S\times T^{\prime}$ are measurable and thus their intersection $S^{\prime}\times T^{\prime}$ is. Since $S\times T$ has a countable basis, every open set is a countable union of open sets and hence $\mathscr{B}$ generates $\mathcal{B}(S\times T)$. Therefore, we get ${\mathcal{B}}(S\times T)\subseteq{\mathcal{B}}(S)\otimes{\mathcal{B}}(T)$. ∎
Also the following is considered to be known; again, we provide a proof for the convenience of the reader.
@@
Let $C_{1},C_{2},C_{3}$ be countable sets. The composition map $\circ\colon C_{3}^{C_{2}}\times C_{2}^{C_{1}}\to C_{3}^{C_{1}}$ is (Borel-)measurable.